Determine the amount of current needed by straight wire 60.0 cm long to generate a 70N force while being exposed on a 27.5 T magnetic field acting 90.0^(0) on the wire.


1
Expert's answer
2022-05-20T13:19:47-0400
"F=BIl\sin\theta"

"I=\frac{F}{Bl\sin\theta}\\=\frac{70.0}{27.5*0.600*\sin90^\circ}=4.24\:\rm A"